(D − 1)^2(D^2 + 1)^2 y = sin^2(x/2) + e^x + x


1
Expert's answer
2021-06-10T07:13:26-0400

Solution

For homogeneous solution let’s solve characteristic equation

(k-1)2(k2+1)=0 => k1,2=1, k3,4=i, k5,6=-i

So the homogeneous solution is yH(x) = (A+Bx)ex+(C+Dx)eix +(E+Fx)e-ix, where A,B,C,D,E,F are arbitrary constants. 

sin2(x/2) + ex + x = x+1/2-cos(x)/2+ex = x+1/2 - eix /4 - e-ix /4 +ex    

Four parts of partial solutions

y1(x) =ax+b: (D − 1)2(D2 + 1)2y1(x) = (D − 1)2(D2 + 1)(ax+b) = (D − 1)2 (ax+b) = (D − 1)(a-ax-b) = -2a+ax+b

ax+b-2a = x+1/2 =>a=1, b=5/2 => y1(x) =x+5/2

y2(x) =cx2ex: (D − 1)2(D2 + 1)2y2(x) = (D2 + 1)2(D − 1)2 cx2ex = c(D2 + 1)2(D − 1) (2xex+x2ex- x2ex) = c(D2 + 1)2(D − 1) (2xex) = c(D2 + 1)2 (2ex+2xex-2xex) = 2c(D2 + 1)2 ex = 2c(D2 + 1) (2ex ) = 8cex => c=1/8 +> y2(x) =x2ex /8

y3(x) =dx2eix: (D − 1)2(D2 + 1)2y3(x) = d(D − 1)2(D2 + 1) ( 2eix + 2ixeix +2ixeix - x2eix + x2eix) = d(D − 1)2 (D2 + 1) (2eix + 4ixeix) = d(D − 1)2 (-2eix - 4eix - 4eix – 4ixeix + 2eix + 4ixeix) = d(D − 1)2 ( - 4eix - 4eix) = -8d(D − 1)2 (eix ) = -8d(D − 1) (ieix - eix ) = -8d(-eix - ieix - ieix + eix ) = 16d ieix => 16d i =-1/4 => d= i /64 => y3(x) = ix2eix/64

y4(x) =fx2e-ix: (D − 1)2(D2 + 1)2y4(x) = f(D − 1)2(D2 + 1) (2e-ix – 2ixe-ix - 2ixe-ix - x2e-ix + x2e-ix) = f(D − 1)2 (D2 + 1) (2e-ix – 4ixe-ix) = f(D − 1)2 (-2e-ix – 4e-ix – 4e-ix + 4ixe-ix +2e-ix – 4ixe-ix) = f(D − 1)2 ( - 4e-ix - 4e-ix) = -8f(D − 1)2 (e-ix ) = -8f(D − 1) (-ie-ix - e-ix ) = -8f(-e-ix + ie-ix + ie-ix + e-ix ) = -16f ie-ix => -16f i =-1/4 => f= -i /64 => y4(x) =-ix2e-ix/64

Solution of the given equation:

y(x) = yH(x) + y1(x) + y2(x) + y3(x)+ y4(x) = (A+Bx)ex+(C+Dx)eix +(E+Fx)e-ix + x+5/2 + x2ex /8 + ix2eix/64 - ix2eix/64 = (A+Bx)ex+(C+Dx)eix +(E+Fx)e-ix + x+5/2 + x2ex/8 - x2sin(x)/32

Answer

y(x) = (A+Bx)ex+(C+Dx)eix +(E+Fx)e-ix + x+5/2 + x2ex/8 - x2sin(x)/32